This year, serena is one-seventh of her cousin's age. Two years from now, serena will be one-fourth of her cousein's age. In how many years will serena be half her cousin's age?
Answer by Alan3354(69443)

This year, serena is one-seventh of her cousin's age. Two years from now, serena will be one-fourth of her cousein's age. In how many years will serena be half her cousin's age?
-----------
S = C/7
S+2 = (C+2)/4
Find their ages.
Sub for S in the 2nd eqn:
C/7 + 2 = (C+2)/4
4C + 56 = 7(C+2) = 7C+14
C = 14
S = 2
===============
Difference = 12 years.
--> 1/2 at ages 12 & 24
----
In how many years will serena be half her cousin's age?
In 10 years.